OpenStudy (anonymous):

The sum of two numbers is 19. Their product is 84. Find the numbers

OpenStudy (anonymous):

12

OpenStudy (anonymous):

so the sum is 19 find all of the numbers that ADD up to 19 1+ 18, 2 + 17, 3 + 16, 4 + 15, 5 +14, 6 + 13, 7 + 12, 8 + 11, 9 + 10 okay, so the product is 84 out of those numbers, which one will have a product of 84? 1 * 18 = 18 2 * 17 = 34 3 * 16 = 48 4 * 15 = 60 5 * 14 = 70 6 * 13 = 78 7 * 12 = 84

OpenStudy (acxbox22):

\(\Huge\color{orange}{\bigstar}\huge\color{blue}{WELCOME~TO~OPENSTUDY}\Huge\color{orange}{\bigstar}\)

OpenStudy (acxbox22):

using @jadeybells chart the last pair of numbers 7 and 12 work there is a proper algebraic way to do it, but her method should work for now

OpenStudy (anonymous):

I do it in a way simpler way and it still comes out with the right answer. Lol no wonder math is so frustrating in school, people always take the long way about things.

OpenStudy (acxbox22):

@jadeybells what would you do if your problem was 2 number that add up to 30987 and mulitply to 19087

OpenStudy (anonymous):

Lol... When that problem actually comes around i'll use the 'proper' way

OpenStudy (acxbox22):

you would have to use algebra \(\Huge\color{orange}{\ddot\smile}\Huge\scr\color{maroon}{LOL}\)

OpenStudy (anonymous):

Yeah... Algebra :( xD I can do it too xD